OpenStudy (anonymous):

6x+ky=10 , m=2 find value of k so given line has slope of m

OpenStudy (anonymous):

-3

OpenStudy (jhannybean):

Solve for y by isolating it on one side and everything else on the other. Subtract -6x from both sides. -6x +6x + ky =10 - 6x Simplify ky = -6x+10 Divide by k both sides of the equation y = -6x/k +10/k Now solve for k when m = 2 -6/k = 2 2k = -6 k= -3 So your equation becomes y = -6x/-3 + 10/-3 y = 2x - 10/3

OpenStudy (jhannybean):

Understand?